Question 38
ProbabilityA box contains 4 red and 3 blue identical balls. If two balls are picked at random, one after the other without replacement, find the probability that one is red and the other is blue.
- A.4/7
- B.2/7
- C.1/7
- D.1/12
Show the answer and explanation
Answer: A (4/7)
Explanation
P(red then blue) + P(blue then red) = (4/7)(3/6) + (3/7)(4/6) = 2/7 + 2/7 = 4/7.